# Frequency shifts as a reflection of ground state squeezing and entanglement in two coupled harmonic oscillators

### v1

@@ Abstract
It is often argued that two coupled quantum harmonic oscillators, even when cooled to their ground state, display no inherently quantum features beyond quantized energy levels. Here, we challenge this view by showing that their classical observables can encode genuinely quantum features. In particular, we demonstrate that the characteristic frequency shifts observed in such systems act as a signature of non-classical correlations and ground-state entanglement at zero temperature, specifically two-mode squeezing between the uncoupled modes. From a complementary perspective, these two effects—frequency shifts and squeezing—represent the same underlying phenomenon expressed in different mode bases. What appears as a spectral renormalization in one description manifests as entanglement in another. These shifts therefore can serve as an entanglement witness accessible via standard frequency measurements. Furthermore, we show that this underlying squeezing, although not directly measurable, can be exploited to enhance the signal-to-noise ratio in precision frequency measurements of individual oscillators without requiring squeezed quantum noise. Our results uncover a new route to quantum-enhanced sensing within systems traditionally regarded as classical, offering fresh insight into how signatures of quantumness persist across the quantum-to-classical boundary.
I Introduction
@@
Coupled harmonic oscillators are among the most fundamental and versatile systems in physics [PhysRev.175.286, Kim2005, PhysRevE.97.042203]. They serve as minimal models for a wide range of physical phenomena [Kim2007], from vibrational modes in molecules [Wilson1955, Ikeda1999, Fillaux2005, Delor2017] and phonons in solids [Girvin2019] through biophysics [Halpin2014, Fuller2014, Romero2014] to cavity modes in quantum optics [Law1994, Walther2006, aspelmeyer2014, PhysRevA.110.063703] and resonators in circuit quantum electrodynamics [Blais2021]. Due to their analytic tractability and ubiquity [oriol2019quadratic], they have played a central role in developing both classical and quantum theories of motion, decoherence [Caldeira1983, Feynman2000, Weiss2011], and measurement [Braginsky1992, Gardiner2004, Wiseman2009]. When two quantum harmonic oscillators are coupled linearly, the resulting system remains Gaussian and exactly solvable [Adesso2007, Weedbrook2012, Serafini2017, HO_2025_quant], leading to the widespread assumption that such dynamics are, in essence, classical [Walls1994, Eisert2003, Scully1997]. Indeed, many aspects of their behavior—such as normal mode formation, energy exchange, and resonance shifts—can be described entirely within the framework of classical physics [MarionThornton]. As a result, coupled oscillator systems are often viewed as belonging to the classical domain unless explicitly driven into non-Gaussian regimes or subjected to strong measurements that reveal nonlocal correlations. This perception has contributed to the prevailing notion that linearly coupled quantum oscillators are quantum in name only, lacking the genuinely non-classical features typically associated with entanglement or squeezing. Consequently, they are frequently used as benchmarks or pedagogical examples where quantum and classical predictions coincide [Feynman1965, Breuer2002], and where the quantum nature of the system is considered irrelevant for practical purposes.
@@
In this work, we challenge this prevailing view by showing that coupled quantum harmonic oscillators, even when appearing classical in their dynamical behavior, can harbor deeply quantum correlations. Specifically, we reveal that the frequency shifts arising from the coupling can serve as indirect signatures of two-mode squeezing and entanglement between the original uncoupled modes. These correlations, though not directly measurable in standard normal mode observables, leave a classical imprint on the system’s response, allowing the frequency shifts to act as entanglement witnesses. Furthermore, we show that this latent squeezing can be harnessed to enhance the precision of frequency estimation for the individual oscillators. In this way, the system enables quantum-enhanced sensing in a platform typically considered classical. Our results contribute to a more nuanced understanding of the quantum-to-classical transition and point to new opportunities for exploiting quantum resources in minimally engineered or weakly measured systems.
II Two coupled harmonic oscillators
@@
Interacting systems can be coupled to the environment through entangled eigenstates [Ciuti2006inputoutputUSC, blais2011dissipation, Bamba2012, Bamba2013, DeLiberato2014]. In such cases, it becomes impossible to directly measure, detect or destroy the entanglement. To illustrate this, consider two coupled harmonic oscillators described by the Hamiltonian
@@
where $\hat{a}$ ($\hat{b}$) annihilates an excitation with energy $\omega$ ($\Omega$), and $g$ is the coupling strength between the oscillators. The ground state of this system can be written as [HongYi1992, emary2003dickemod, Zhou2020]
@@
where subscripts $c$ and $d$ refer to the eigen bases of now coupled harmonic oscillators with normal frequencies $\omega e^{-2\xi_{-}}$ and $\Omega e^{-2\xi_{+}}$. In other words, while the state may appear classical in a particular basis, the system itself can still be fundamentally entangled. At first glance, this may seem paradoxical: can entanglement really be created or removed by a mere change of basis? The answer is more subtle. To speak meaningfully about entanglement, one must first specify a partition of the system into (at least) two subsystems, and this partition is not unique. A change of basis may correspond to a change in this partitioning, transforming what appears as entanglement in one description into a separable state in another, without altering the physical state itself. This perspective is not only a matter of mathematical representation. In relativistic quantum information it is well known that entanglement is observer dependent. Different observers may decompose the quantum field into different modes, leading to inequivalent partitions of the Hilbert space and hence to different entanglement structures. For instance, uniformly accelerated observers perceive the Minkowski vacuum as a thermal state (the Unruh effect), which degrades the entanglement seen by inertial observers [Alsing2003, Fuentes2005, Peres2004]. More generally, relative motion or spacetime curvature can generate or diminish entanglement depending on the chosen partitioning [Friis2012, MartinMartinez2009].
Measurement in an arbitrary basis is generally nontrivial. Most quantum systems are not measured directly but rather through coupling to an external environment—often via the electromagnetic field [Carmichael1993, Wiseman2009]. For systems with a linear energy structure (such as linearly interacting harmonic oscillators), if driven by a coherent state of light, they will emit a coherent state in return, regardless of the complexity of their internal quantum states [Glauber1963, Gardiner1985, Gardiner2004]. This means that the interacting harmonic oscillators can be coupled to the environment through hybrid entangled eigenmodes described by $\hat{c}$ and $\hat{d}$ rather than $\hat{a}$ and $\hat{b}$. In the language of the master equation, $\hat{c}$ and $\hat{d}$ are jump operators. This suggests that no information about the entanglement of eigenstates is imprinted onto the environment.
@@
[…]
@@
III The inseparability criterion
Let us have a look now at entanglement. In continuous variable quantum systems, such as optical modes described by canonical quadrature operators, entanglement can be detected via the inseparability criterion proposed by Duan et al. [Duan2000] and independently by Simon [Simon2000]. Consider quadrature operators of two bosonic modes, for mode a
@@ IV Finite temperature
The analysis above demonstrates that frequency shifts not only signal entanglement through the inseparability criterion, but also remain quantitatively linked to the degree of two-mode squeezing even at finite temperature. In this sense, the same microscopic correlations that drive the inseparability violation underlie the macroscopic spectral response of the system. This duality suggests a deeper operational meaning of the observed frequency renormalization. It is not merely a static spectral feature, but a manifestation of the very quantum resources that enable enhanced metrological performance. Consequently, the frequency shifts that witness entanglement in equilibrium can also act as carriers of metrological advantage out of equilibrium. In the following, we demonstrate how these hidden correlations—encoded in the system’s structure rather than in its noise statistics—translate into measurable improvements in parameter estimation, revealing a direct bridge between correlations and sensitivity.
V quantum enhanced sensitivity with squeezing the state
To place our results in context, we first revisit the conventional route to quantum-enhanced sensitivity based on explicit squeezing of quantum noise. For this purpose, let us consider a single-mode quantum harmonic oscillator governed by the Hamiltonian
@@
In frequency metrology, one typically prepares a coherent state by displacing the vacuum [Glauber1963, Walls1994, Scully1997] and probes the system’s dynamics to estimate the oscillator frequency [Giovannetti2004, Paris2009, Pezze2018]. A common choice of measurement is the position quadrature, usually implemented via homodyne detection [Leonhardt1997, Gardiner2004].
@@
where $\mathcal{M}$ (in the following set to 1 for simplicity) is the number of measurements performed to construct the expectation value of $\hat{x}$. To improve precision, one can reduce the quantum noise by using a squeezed coherent state of the form
@@
This is precisely why squeezed states are so valuable in quantum metrology. They allow one to surpass the standard quantum limit by amplifying the signal-to-noise ratio through quantum noise reduction.
VI quantum enhanced sensitivity with squeezing the oscillator
Now consider a frequency measurement of an interacting harmonic oscillator from Eq. (6) (we concentrate on one mode only)
@@
while still aiming to estimate the original bare frequency $\omega$ [Gietkacxvs-5pb1]. For a coherent state $|\alpha\rangle_{c}$—squeezed in terms of $\hat{a}$ and $\hat{a}^{\dagger}$ operators—where the subscript $c$ indicates the new basis, it is straightforward to calculate the signal-to-noise ratio. The noise is of course
@@
[…]
this is, the signal is multiplied by a derivative of the effective frequency with respect to the frequency of the non-interacting harmonic oscillator that we want to measure. This derivative can be easily evaluated and for a strongly off-resonant case ($\omega\ll\Omega$ taken for simplicity) turns out to be
@@
Although the system appears entirely classical and no direct squeezing of quantum noise is involved, the signal-to-noise ratio is nonetheless enhanced—governed by the squeezing parameter associated with the inaccessible modes $a$ and $b$. In other words, it is possible to achieve a quantum advantage within a seemingly classical system. This highlights that the classical observables of coupled harmonic oscillators are fundamentally shaped by their underlying quantum correlations. Even in the absence of overt quantum features, quantum resources might remain operative and exploitable. Additionally, one could also create squeezing in the new mode $\hat{c}$ and combine the squeezing effects at two different levels [PhysRevLett.132.060801, mihailescu2025criticalquantumsensingtutorial].
VII conclusions & Discussion
@@
We have demonstrated that linearly coupled quantum harmonic oscillators—systems often regarded as essentially classical due to their Gaussian states and analytical solvability—can harbor hidden quantum correlations that manifest through measurable, seemingly classical observables. In particular, we have shown that coupling-induced frequency shifts encode two-mode squeezing between the original oscillator modes. From a complementary viewpoint, the frequency shifts and the squeezing describe the same physical effect, merely expressed in different mode bases. In the normal-mode picture it appears as a frequency renormalization, while in the original basis it manifests as two-mode squeezing. Although this squeezing might not be directly observable, it leaves a clear imprint on the system’s response, which can be harnessed for quantum-enhanced frequency estimation. Strikingly, this enhancement arises without the need to manipulate quantum noise directly. Instead of relying on traditional noise squeezing to reduce measurement uncertainty, the quantum advantage stems from the implicit squeezing embedded in the ground state of the interacting system. This mechanism leads to a signal-to-noise improvement that scales as $\exp(4\xi)$—an enhancement over the standard $\exp(2\xi)$ scaling achieved through explicit squeezing [Gietkacxvs-5pb1]. This indicates that quantum resources might be more efficiently exploited when encoded into the system’s structure rather than externally imposed.
@@
These findings challenge the commonly held belief that squeezing has no classical counterpart. We propose that the metrological advantage typically attributed to quantum squeezing can, in certain systems, be revealed as classical frequency shifts. This suggests a subtle reinterpretation of squeezing. While the underlying correlations are quantum in origin, their effects may be accessed through classical measurements of system response [Walls1983, Caves1981]. Our results blur the conventional boundary between classical and quantum regimes, showing that systems traditionally considered classical can offer genuine quantum advantages when appropriately interpreted. This sheds new light on the classical-to-quantum transition, suggesting that the emergence of quantum-enhanced behavior need not coincide with abrupt changes in nonclassicality indicators, but may instead unfold gradually—encoded in how a system responds to parameter variations [Kofler2007, Leggett2002]. In this view, classical observables, such as frequencies, can serve as sensitive witnesses of latent quantum correlations, even in regimes where traditional signatures of quantumness are absent [Brunner2014, Adesso2016]. These insights open a promising direction for quantum metrology in platforms where direct quantum control is limited or costly. Weakly interacting or passive systems, which require no state preparation beyond coherent displacements, may still provide access to quantum resources—hidden in plain sight as classical observables [Pezze2018]. Ultimately, this motivates a deeper investigation into how quantum correlations can be embedded, witnessed, and utilized across a broad range of physical systems, reframing our understanding of the quantum–classical divide in both conceptual and practical terms.
@@
The results can be naturally extended to interacting spin systems that admit a mapping to a harmonic oscillator, such as the Lipkin–Meshkov–Glick model [Lipkin1965, LMG2005, LMG2007, LMG2008]. This offers a direct route to certify entanglement—in the form of spin squeezing—by simply monitoring how the energy gap is modified by spin–spin interactions. A more intriguing challenge, however, is to uncover how the energy gap relates to spin squeezing in less symmetric systems that cannot be reduced to a harmonic oscillator description. We anticipate that this connection will hinge on the interplay between the gap and the symmetry properties of the eigenstates [Gietka2025prep], potentially opening novel pathways for generating scalable spin squeezing in many-body spin systems. Establishing such a relation would not only broaden the range of experimentally accessible platforms, but could also provide a unifying framework linking spectral properties and metrological performance in complex quantum matter.
Acknowledgements.

### v2

# Frequency shifts heralding ground state squeezing and entanglement of two coupled harmonic oscillators

@@ Abstract
It is often argued that two linearly coupled quantum harmonic oscillators, even when cooled to their ground state, display no inherently quantum features beyond quantized energy levels. Here, we challenge this view by showing that their classical observables encode genuinely quantum features. In particular, we demonstrate that the characteristic frequency shifts observed in coupled oscillators signal non-classical correlations and ground-state entanglement at zero temperature corresponding to two-mode squeezing between the uncoupled modes. From a complementary perspective, these two effects, frequency shifts and squeezing, represent the same underlying phenomenon but expressed in different mode bases. What appears as a spectral renormalization in one description manifests itself as entanglement in the other. Frequency shifts therefore constitute an entanglement witness accessible via standard spectroscopy. While the underlying squeezing is not directly measurable, it can be exploited to enhance the signal-to-noise ratio in precision frequency measurements of individual oscillators without requiring squeezed quantum noise. This uncovers a new route to quantum-enhanced sensing within systems traditionally regarded as classical, offering fresh insight into how signatures of quantumness persists across the quantum-to-classical boundary.
1 Introduction
@@
Coupled harmonic oscillators are among the most fundamental and commonly used systems in physical modeling [1, 2, 3]. They serve as minimal but solvable models for a wide range of physical phenomena [4], from vibrational modes in molecules [5, 6, 7, 8], phonons in solids [9] through biophysics [10, 11, 12] to cavity modes in quantum optics [13, 14, 15, 16] and circuit quantum electrodynamics [17]. Due to their analytic tractability and ubiquity [18], they played a central role in developing both classical and quantum theories of motion, decoherence [19, 20, 21], and measurement [22, 23, 24]. When two quantum harmonic oscillators are coupled linearly, the resulting system remains Gaussian and exactly solvable [25, 26, 27, 28, 29, 30], leading to the widespread conjecture that this dynamics is, in essence, classical [31, 32, 33]. Indeed, many aspects of the coupled dynamics—such as normal mode formation, energy exchange, and resonance shifts—can be derived entirely from the framework of classical physics [34]. As a result, coupled oscillator systems are commonly viewed as belonging to the classical domain unless explicitly driven into non-Gaussian regimes or subjected to strong measurements that reveal nonlocal correlations. This perception has contributed to the prevailing notion that linearly coupled quantum oscillators are quantum in name only, lacking the genuinely non-classical features as entanglement or squeezing. Hence they are frequently used as benchmarks or pedagogical examples where quantum and classical predictions coincide [35, 36] and the quantum nature of the system isirrelevant.
@@
In this work, we challenge this prevailing view by showing that coupled quantum harmonic oscillators, even when appearing classical in their dynamical behavior, can harbor deeply quantum correlations. Specifically, we reveal that the frequency shifts arising from the coupling can serve as indirect signatures of two-mode squeezing and entanglement between the original uncoupled modes. These correlations, though not directly measurable in standard normal mode observables, leave a classical imprint on the system’s response, allowing the frequency shifts to be used as entanglement witnesses. Furthermore, we show that the inherent latent squeezing can be harnessed to enhance the precision of frequency estimation for the individual oscillators. In this way, the system enables quantum-enhanced sensing in a platform typically considered classical. Our results contribute to a more nuanced understanding of the quantum-to-classical transition and point to new opportunities for exploiting quantum resources in minimally engineered or weakly measured systems.
2 Two coupled harmonic oscillators
@@
Interacting systems can be coupled to the environment through entangled eigenstates [37, 38, 39, 40, 41]. In such cases, it becomes impossible to directly measure, detect or destroy the entanglement. To illustrate this, consider two coupled harmonic oscillators described by the Hamiltonian
@@
where $\hat{a}$ ($\hat{b}$) annihilates an excitation with energy $\omega$ ($\Omega$), and $g$ is the coupling strength between the oscillators. The ground state of this system can be written as [42, 43, 44]
@@
where subscripts $c$ and $d$ refer to the eigen bases of now coupled harmonic oscillators with normal frequencies $\omega e^{-2\xi_{-}}$ and $\Omega e^{-2\xi_{+}}$. In other words, while the state may appear classical in a particular basis, the system itself can still be fundamentally entangled. At first glance, this may seem paradoxical: can entanglement really be created or removed by a mere change of basis? The answer is more subtle. To speak meaningfully about entanglement, one must first specify a partition of the system into (at least) two subsystems, and this partition is not unique. A change of basis may correspond to a change in this partitioning, transforming what appears as entanglement in one description into a separable state in another, without altering the physical state itself. This perspective is not only a matter of mathematical representation. In relativistic quantum information it is well known that entanglement is observer dependent. Different observers may decompose the quantum field into different modes, leading to inequivalent partitions of the Hilbert space and hence to different entanglement structures. For instance, uniformly accelerated observers perceive the Minkowski vacuum as a thermal state (the Unruh effect), which degrades the entanglement seen by inertial observers [45, 46, 47]. More generally, relative motion or spacetime curvature can generate or diminish entanglement depending on the chosen partitioning [48, 49].
Measurement in an arbitrary basis is generally nontrivial. Most quantum systems are not measured directly but rather through coupling to an external environment—often via the electromagnetic field [50, 24]. For systems with a linear energy structure (such as linearly interacting harmonic oscillators), if driven by a coherent state of light, they will emit a coherent state in return, regardless of the complexity of their internal quantum states [51, 52, 23]. This means that the interacting harmonic oscillators can be coupled to the environment through hybrid entangled eigenmodes described by $\hat{c}$ and $\hat{d}$ rather than $\hat{a}$ and $\hat{b}$. In the language of the master equation, $\hat{c}$ and $\hat{d}$ are jump operators. This suggests that no information about the entanglement of eigenstates is imprinted onto the environment.
@@
$\xi_{-}$ $=-\frac{1}{2}\log{\frac{\omega_{-}}{\omega}},$ (8)
$\xi_{+}$ $=-\frac{1}{2}\log{\frac{\omega_{+}}{\Omega}},$ (9)
@@
3 The inseparability criterion
Let us have a look now at entanglement. In continuous variable quantum systems, such as optical modes described by canonical quadrature operators, entanglement can be detected via the inseparability criterion proposed by Duan et al. [53] and independently by Simon [54]. Consider quadrature operators of two bosonic modes, for mode $a$
@@ 4 Finite temperature
The analysis above demonstrates that frequency shifts not only signal entanglement through the inseparability criterion, but also remain quantitatively linked to the degree of two-mode squeezing even at finite temperature. Note that the same microscopic correlations that induce inseparability underlie the macroscopic spectral response of the system. This duality suggests a deeper operational meaning of the observed frequency renormalization. It is not merely a static spectral feature, but a manifestation of the very quantum resources that enable enhanced metrological performance. Consequently, the frequency shifts that witness the entanglement will also carry the metrological advantage. In the following we demonstrate how these hidden correlations—encoded in the system’s structure rather than in its noise statistics—translate into measurable improvements in parameter estimation, revealing a direct bridge between correlations, sensitivity, and frequency shifts.
5 Quantum enhanced sensitivity through squeezed states
Let us first revisit the conventional route to quantum-enhanced sensitivity based on squeezing of quantum noise. For this purpose we consider a single quantum harmonic oscillator governed by the Hamiltonian
@@ 5 Quantum enhanced sensitivity through squeezed states
In frequency metrology, one typically prepares a coherent state by displacing the vacuum [51, 31, 33] and monitors the system’s dynamics to estimate the oscillator frequency [55, 56, 57]. A common choice of measurement is the position quadrature implemented via homodyne detection [58, 23].
@@
where $\mathcal{M}$ (in the following set to 1 for simplicity) is the number of measurements performed to construct the expectation value of $\hat{x}$. To improve precision one reduces the quantum noise by using a squeezed coherent state of the form
@@
This is precisely why squeezed states are so valuable in quantum metrology. They allow one to surpass the standard quantum limit by enlarging the signal-to-noise ratio through quantum noise reduction.
6 Quantum enhanced sensitivity with squeezing the oscillator’s frequency
Now consider a frequency measurement on an interacting harmonic oscillator from Eq. (6) (here we concentrate on one of the modes only)
@@ 6 Quantum enhanced sensitivity with squeezing the oscillator’s frequency
still with the aim to precisely estimate the bare frequency of this oscillator without interaction $\omega$ [59]. For a coherent state $|\alpha\rangle_{c}$—squeezed in terms of $\hat{a}$ and $\hat{a}^{\dagger}$ operators—where the subscript $c$ indicates the new basis, it is straightforward to calculate the signal-to-noise ratio. The noise is of course
@@
$\partial_{\omega}\langle\hat{x}_{c}\rangle=|\alpha|{t}\cos{\omega_{-}t}\times\partial_{\omega}\omega_{-}.$ (30)
We find that the signal is multiplied by a derivative of the effective frequency with respect to the frequency of the non-interacting harmonic oscillator that we want to measure. This derivative can be explicitly evaluated. For a sufficiently far off-resonant case ($\omega\ll\Omega$ taken for simplicity) it turns out to be
@@
Although the system appears entirely classical and no direct squeezing of quantum noise is involved, the signal-to-noise ratio is nonetheless enhanced. The improvement is governed by the squeezing parameter associated with the inaccessible modes $a$ and $b$. In other words, it is possible to obtain a "quantum" enhanced measurement within a seemingly classical system. This highlights that the classical observables of coupled harmonic oscillators are fundamentally shaped by their underlying quantum correlations. Even in the absence of overt quantum features, quantum resources might remain operative and exploitable. Additionally, one could also create squeezing in the new mode $\hat{c}$ and combine the squeezing effects at two different levels [60, 61].
7 Conclusions & discussion
@@
We have demonstrated that linearly coupled quantum harmonic oscillators—systems, often regarded as essentially classical due to their Gaussian states and analytical solvability, can harbor hidden quantum correlations. These manifest themselves through measurable, seemingly classical observables. In particular, we have shown that coupling-induced frequency shifts encode two-mode squeezing between the original oscillator modes. From two complementary viewpoints, the frequency shifts and the squeezing describe the same physical effect of coupling, merely expressed in different mode bases. In the normal-mode picture it appears as a frequency renormalization, while in the original basis it manifests as two-mode squeezing.
Although the underlying squeezing might not be directly observable, it leaves a clear imprint on the system’s response, which can be harnessed for quantum-enhanced frequency estimation. Strikingly, this enhancement arises without the need to manipulate quantum noise directly. Instead of relying on traditional noise squeezing to reduce measurement uncertainty, the quantum advantage stems from the implicit squeezing embedded in the eigenstates of the interacting system. This mechanism leads to a signal-to-noise improvement that scales as $\exp(4\xi)$—an enhancement over the standard $\exp(2\xi)$ scaling achieved through explicit squeezing [59]. Hence in suitable cases, quantum resources might be more efficiently exploited when encoded into the system’s structure rather than externally imposed.
Our findings challenge the common held belief that squeezing has no classical counterpart. We propose that the metrological advantage typically attributed to quantum squeezing can in certain systems appear as classical frequency shifts. This suggests a subtle reinterpretation of squeezing. While the underlying correlations are quantum in origin, their effects may be accessed through classical measurements of the system response [62, 63]. Our results blur the conventional boundary between classical and quantum regimes, showing that systems traditionally considered classical can allow for equal measurement advantages as a quantum model when appropriately interpreted.
@@
From a wider perspective this also sheds some new light on the classical-to-quantum transition, suggesting that the emergence of quantum-enhanced behavior need not coincide with abrupt changes in nonclassicality indicators; it may instead unfold gradually encoded in how a system responds to parameter variations [64, 65]. In this view, classical observables, such as frequencies, can serve as sensitive witnesses of latent quantum correlations, even in regimes where traditional signatures of quantum-ness are absent [66, 67]. These opens a promising direction for quantum metrology in platforms where direct quantum state control is limited. Weakly interacting or passive systems, which require no state preparation beyond coherent displacements, may still provide access to quantum resources—hidden in plain sight as classical observables [57]. Ultimately, this motivates a deeper investigation into how quantum correlations can be embedded, witnessed and utilized across a broad range of physical systems, reframing our understanding of the quantum–classical divide in both conceptual and practical terms.
@@
Let us finally remark, that the results can be naturally extended to interacting spin systems that admit a mapping to a harmonic oscillator, such as the Lipkin–Meshkov–Glick model [68, 69, 70, 71]. This offers a direct route to certify entanglement—in the form of spin squeezing—by simply monitoring how the energy gap is modified by spin–spin interactions. A more intriguing challenge, however, is to uncover how the energy gap relates to spin squeezing in less symmetric systems that cannot be reduced to a harmonic oscillator description. We anticipate that this connection will hinge on the interplay between the gap and the symmetry properties of the eigenstates [72], potentially opening novel pathways for generating scalable spin squeezing in many-body spin systems. Establishing such a relation would not only broaden the range of experimentally accessible platforms, but could also provide a unifying framework linking spectral properties and metrological performance in complex quantum matter.
Acknowledgements